A shop has three legged stools and 4 legged stools. There are 31 stools and 104 legs. How many of each type of stools are there?
Answer by cleomenius(959)

x + y = 31
3x + 4y = 104
-3x + - 3y = -93
3x + 4y = 104
y = 11
x = 20
20(3) + 11(4) = 104
Cleomenius.